OpenStudy (anonymous):

find range of f(x)=(x-2)/(2-x)

OpenStudy (mathmale):

Note that your f(x)=(x-2)/(2-x) is equivalent to\[f(x)=-(x-2)/(x-2)\] Does that tell you anything about the domain and range of this function?

OpenStudy (anonymous):

so....-1 is the answer?

OpenStudy (anonymous):

@mathmale ..-1 is the answer?

OpenStudy (mathmale):

\[f(x)=-(x-2)/(x-2)=-1\] is true. But is -1 the answer? No. What are you asked to do in this problem?

OpenStudy (anonymous):

to find the range..

OpenStudy (anonymous):

means value of f(x)

OpenStudy (anonymous):

hmm....yess...it is -1..:)

OpenStudy (anonymous):

thanks for the hint

OpenStudy (mathmale):

Actually, the "range" of a function is the set of all POSSIBLE function values (outcomes). Does it seem reasonable to say that the range of this function is just {-1}?

OpenStudy (anonymous):

yes...giving any value to x gives us -1

OpenStudy (mathmale):

Nice work! Take care!